# Time-independent Spiking Neuron via Membrane Potential Estimation for Efficient Spiking Neural Networks

### v1

@@ Abstract
The computational inefficiency of spiking neural networks (SNNs) is primarily due to the sequential updates of membrane potential, which becomes more pronounced during extended encoding periods compared to artificial neural networks (ANNs). This highlights the need to parallelize SNN computations effectively to leverage available hardware parallelism. To address this, we propose Membrane Potential Estimation Parallel Spiking Neurons (MPE-PSN), a parallel computation method for spiking neurons that enhances computational efficiency by enabling parallel processing while preserving the intrinsic dynamic characteristics of SNNs. Our approach exhibits promise for enhancing computational efficiency, particularly under conditions of elevated neuron density. Empirical experiments demonstrate that our method achieves state-of-the-art (SOTA) accuracy and efficiency on neuromorphic datasets without requiring additional training parameters. Codes are available at https://github.com/chrazqee/MPE-PSN.
@@ 1 Introduction
Fig. 1: Variations in the execution time ratio between the vanilla LIF neuron and the MPE-PSN neuron concerning time step increments and neuron quantity augmentation in both cases during the forward process, denoted as $\frac{T_{vanilla\leavevmode\nobreak\ LIF}}{T_{MPE-PSN}}$.
To address prevailing challenges, we introduce a Membrane Potential Estimation Parallel Spiking Neuron (MPE-PSN), a highly dynamic and fully parallelized computational neuron. MPE-PSN estimates the activation output through the spiking probability of the membrane potential , thereby decoupling spiking neuron computations from temporal dependencies. This strategy promises a substantial boost in the computational efficiency of long-temporal dependent spiking neural networks, exemplified by [7]. Throughout SNN training, the alignment between true and estimated membrane potentials is optimized using the minimum mean square error (MSE) loss function. Empirical findings validate the marked improvement in computational efficiency offered by the proposed methodology, while preserving the intricate dynamic attributes of spiking neurons. Noteworthy performance benchmarks are attained on neuromorphic datasets.
@@ 2 Methods
Fig. 2: Schematic diagram of the proposed neuron internal structure based on membrane potential estimation.
@@ 2.1 Motivation and Derivation of Neuron Design
Spiking neurons excel in temporal information processing. Their iterative computational process can be outlined as follows:
@@
$o_{i,t}^{\ell}$ $=\Theta(h_{i,t}^{\ell}-v_{th})=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \leavevmode\nobreak\ u_{i,t}^{\ell}\geq v_{th},\\ 0,\leavevmode\nobreak\ \leavevmode\nobreak\ otherwise,\end{array}\right.$ (4)
@@
where $I_{i,t}^{\ell}$ denotes the input current of the $i$-th neuron in the $l$-th layer at timestep $t$, $u_{i,t}^{\ell}$ represents the membrane potential post-spike firing and reset, and $h_{i,t}^{\ell}$ signifies the integrated membrane potential prior to firing. The threshold for membrane potential is denoted as $v_{th}$, and $o_{i,t}^{\ell}$ represents the output following spike firing. The function $g(\cdot)$ in Equation (1) corresponds to distinct spiking neurons, while $\Theta(\cdot)$ in Equation (4) signifies the Heaviside step function. In Equation (5), $v_{r}$ stands for the reset potential. This equation elucidates a hard reset scenario in a spiking neuron, where the membrane potential is reset from $u_{i,t}^{\ell}$ to $v_{r}$ upon firing. In surrogate training, the hard reset method is favored over the soft reset approach for enhanced performance [8].
By extending the Equations (1) to (5) to Leaky Integrate-and-Fire (LIF) neuron, we obtain the Equation (6) as:
$u_{i,t}^{\ell}=(\tau_{m}u_{i,t-1}^{\ell}+I_{i,t}^{\ell})(1-\Theta(\tau_{m}u_{i,t-1}^{\ell}+I_{i,t}^{\ell},v_{th}),$ (6)
where $\tau_{m}$ represents the membrane potential constant, input decay is disregarded, and $v_{r}$ is set to 0 in this work, the iterative update of a spiking neuron adheres to a sequential, incremental process. At time step $t$, the output is contingent on the preceding time step $t-1$. The computational time complexity of the neuron at this stage is $\mathcal{O}(T)$, with $T$ denoting the total number of time steps. Given the assumption of independence of membrane potential updates from the previous time point’s output, spiking neuron updates can be architected for parallel processing.
Apparently, when $t=0$, the previous time point’s membrane potential of the neuron is noted as 0, $u_{i,-1}^{\ell}=0$. Consequently, the computation for $u_{i,0}^{\ell}$ is exact. Referring to Equation 6, we obtain:
$u_{i,0}^{\ell}=I_{i,0}^{\ell}(1-\Theta(I_{i,0}^{\ell},v_{th})),$ (7)
@@
$I_{i,t}^{l}$ $=\sum_{j=1}^{N_{l-1}}w_{i,j}^{\ell-1}o_{j,t-1}^{\ell-1},$ (8)
Here, $w_{i,j}^{\ell-1}$ represents the synaptic weight between the $j$-th neuron in the preceding layer and the $i$-th neuron in the subsequent layer. $N_{l-1}$ signifies the number of neurons in the $(l-1)$-th layer.
The accurate $u_{i,0}^{\ell}$ is utilized to rectify the error, denoted as $\mathcal{E}$. The process unfolds as follows:
$u_{i,t}^{\ell}$ $=(\tau_{m}\hat{u}_{i,t-1}^{\ell}+I_{i,t}^{\ell})(1-\Theta(\tau_{m}\hat{u}_{i,t-1}^{\ell}+I_{i,t}^{\ell},v_{th}),$ (9)
$u_{i,t+1}^{\ell}$ $=(\tau_{m}\hat{u}_{i,t}^{\ell}+I_{i,t+1}^{\ell})(1-\Theta(\tau_{m}\hat{u}_{i,t}^{\ell}+I_{i,t+1}^{\ell},v_{th}),$ (10)
@@ 2.2 Initialization of Estimated Membrane Potential
The output of spiking neurons, denoted by the tensor $\mathbf{o}\in\mathbb{R}^{T\times C\times H\times W}$, with $T$, $C$, $H$, and $W$ representing time step, channel, height, and width respectively, adheres to the following distribution:
$\mathbf{o}\sim\mathcal{N}(p,p(1-p)),\leavevmode\nobreak\ \leavevmode\nobreak\ i.e.\leavevmode\nobreak\ Bernoulli(),\leavevmode\nobreak\ \mathbf{o}\in\{0,1\}$ (12)
The output characteristics of individual spiking neurons adhere to a $Bernoulli$ distribution, where $p$ and $1-p$ represent the probabilities of 1 and 0, respectively. Activated neurons produce spikes, while silent neurons do not. Spike generation is influenced by the input current magnitude, with neurons receiving stronger input currents having a higher likelihood of generating spikes. As outlined in [9], this can be described as:
$\mathbf{o}\sim Bernoulli(\mathbb{P}[o_{i,t}^{\ell}=1]),$ (13)
Here, $\mathbb{P}[o_{i,t}^{\ell}=1]=\mathrm{F}_{\epsilon}(\cdot)$, where $\mathrm{F}_{\epsilon}[\cdot]$ denotes the cumulative distribution function of the input current $\mathbf{I}$ in this context, and $\mathbb{P}[\cdot]$ signifies probability.
In our study, we employ the $sigmoid$ function as the $\mathrm{F}_{\epsilon}[\cdot]$, utilized for evaluating the firing probabilities of each neuron.
$\mathbb{P}[\mathbf{o}]$ $=sigmoid(\mathbf{I}),\leavevmode\nobreak\ \mathbf{I}\in\mathbb{R}^{T\times C\times H\times W}$ (14)
$\mathbf{\hat{u}}$ $=(1-Bernulli(\mathbb{P}[\mathbf{o}]))\odot\mathbf{I}.$ (15)
@@
Here, the symbol $\odot$ represents the Hadamard product. The time complexity for computing $\mathbb{P}[\mathbf{o}]$ is theoretically $O(1)$, similar to the $Bernoulli$ operator. As a result, the extra time needed for this operation is negligible, especially on graphics processing unit (GPU) devices.
@@ 2.3 Membrane Potential Approximation Loss
$\mathcal{L}_{mem}=\sum_{t=0}^{T-1}\sum_{i=0}^{N_{l}-1}\kappa_{i}MSE(\hat{u}_{i,t}^{\ell},{u}_{i,t}^{\ell}),$ (16)
Conclusively, the overall loss is calculated as:
@@
where $\lambda$ serves as a coefficient, set to 0.01. $\mathcal{L}_{\text{cls}}$ denotes the classification loss, which is computed employing the methodology outlined in the [12]. $\mathbf{o}_{t}$ and $\mathbf{y}$ represent outputs and labels, respectively. The loss function of $\mathcal{L}_{cls}$ is:
$\mathcal{L}_{cls}=\frac{1}{T}\sum_{t=0}^{T-1}\mathcal{L}_{CE}[\mathbf{o}_{t},\mathbf{y}].$ (18)
@@ 3 Expriments
$\frac{\partial L}{\partial w_{i,j}^{\ell}}=\sum_{t=1}^{T}\frac{\partial L}{\partial o_{i,t}^{\ell}}\frac{\partial o_{i,t}^{\ell}}{\partial{u}_{i,t}^{l}}\frac{\partial{u}_{i,t}^{\ell}}{\partial I_{i,t}^{\ell}}\frac{\partial I_{i,t}^{\ell}}{\partial w_{i,j}^{\ell}},$ (19)
@@
$\frac{\partial o_{i,t}^{\ell}}{\partial{u}_{i,t}^{l}}=\frac{1}{\alpha^{2}}max(0,\alpha-|{u}_{i,t}^{l}-v_{th}|).$ (20)
here, $\alpha$ serves as a constant that restricts the range of activated gradients, fixed at 1.0. The hyper-parameter configurations and network architecture details are outlined in Table 1. Additionally, $\tau_{m}$ is set to 0.25, with the threshold represented as a parameter subject to learning.
@@ 3.1 Comparisons With Existing Methods
DVSGesture [14]. We performed extensive classification experiments on the DVSGesture dataset using the 7B-Net architecture. Our methodology achieved a top-1 accuracy of 97.92% at 16 time steps. This outcome surpassed the SOTA result with 4 fewer time steps.
Table 2: Compare with previous studies on neuromorphic and static datasets. T represents time-steps, while B indicates learnable blocks. Prefix M signifies the term modified. Results marked with $\ast$ denote self-implemented findings. The threshold parameter for Vanilla LIF and PLIF is established at 1.0 for the N-Caltech 101 dataset.
@@
DVS- Gesture KLIF  [4] M-PLIF-Net 12 94.1
@@
MPE-PSN 7B-Net 16 97.92
CIFAR10- DVS GLIF [3] 7B-wideNet 16 78.1
@@
N-Caltech 101 [18]. Our method’s efficacy was further confirmed on the N-Caltech 101 dataset, with the performance of other methods reproduced for comparison. The outcomes revealed that our approach achieved the highest accuracy presently attainable on the N-Caltech 101 dataset.
@@ 3.2 Membrane Potential Approximation Analysis
(a) $L_{2}\leavevmode\nobreak\ norm(\hat{\mathbf{u}},\mathbf{u})$ and Accuracy
@@
(c) Acc. and Loss curve
Fig. 4: 2D loss contour [20] as well as accuracy and loss curve on CIFAR10DVS dataset when time steps is 4.
@@ 3.3 Computational Efficiency Analysis
A comparative analysis is conducted to assess the computational speed difference between our proposed neuron model and the traditional LIF neuron across various time steps and neuron quantities. As depicted in Fig. 1, our neurons generally outpace vanilla LIF neurons in simulation speed. Furthermore, as shown in Fig. 3 (b), our proposed neuron exhibits a notably higher spike rate across most layers when utilizing probabilistic estimation, contrasting with the dynamic complexity observed in the vanilla LIF neuron. Fig. 4 (a) and 4 (b) illustrate that the loss contours of our proposed neurons are comparable to those of vanilla LIF neurons, with both displaying relatively flat profiles. Fig. 3 (a) and 4 (c) demonstrate that initial uncertainties in membrane potential estimation did impact the network’s fitting performance. However, as the estimation of membrane potential improved, this effect became less pronounced. Consequently, our proposed neurons maintain both computational efficiency and performance integrity.

### v2

@@ Abstract
The computational inefficiency of spiking neural networks (SNNs) is primarily due to the sequential updates of membrane potential, which becomes more pronounced during extended encoding periods compared to artificial neural networks (ANNs). This highlights the need to parallelize SNN computations effectively to leverage available hardware parallelism. To address this, we propose Membrane Potential Estimation Parallel Spiking Neurons (MPE-PSN), a parallel computation method for spiking neurons that enhances computational efficiency by enabling parallel processing while preserving the intrinsic dynamic characteristics of SNNs. Our approach exhibits promise for enhancing computational efficiency, particularly under conditions of elevated neuron density. Empirical experiments demonstrate that our method achieves state-of-the-art (SOTA) accuracy and efficiency on neuromorphic datasets. Codes are available at https://github.com/chrazqee/MPE-PSN.
@@ 1 Introduction
Fig. 1: Variations in the execution time ratio between the vanilla-LIF neuron and the MPE-PSN neuron concerning time step increments and neuron quantity augmentation in both cases during the forward process, denoted as $\frac{T_{vanilla-LIF}}{T_{MPE-PSN}}$.
To address prevailing challenges, we introduce a Membrane Potential Estimation Parallel Spiking Neuron (MPE-PSN), a highly dynamic and fully parallelized computational neuron. MPE-PSN estimates the activation output through the spiking probability of the membrane potential , thereby decoupling spiking neuron computations from temporal dependencies. This strategy promises a substantial boost in the computational efficiency of long-temporal dependent spiking neural networks, as exemplified by [7]. Throughout SNN training, the alignment between true and estimated membrane potentials is optimized using the minimum mean square error (MSE) loss function. The empirical findings validate the marked improvement in computational efficiency offered by the proposed methodology while preserving the intricate dynamic attributes of spiking neurons. Notable performance benchmarks are attained on neuromorphic datasets.
@@ 2 Methods
Fig. 2: Schematic diagram of the proposed spiking neuron internal structure based on membrane potential estimation.
@@ 2.1 Motivation and Derivation of Neuron Design
Spiking neurons perform efficiently in temporal information processing. Their iterative computational process can be outlined as follows:
@@
$o_{i,t}^{\ell}$ $=\Theta(h_{i,t}^{\ell}-v_{th})=\left\{\begin{array}[]{l}1,\leavevmode\nobreak\ \leavevmode\nobreak\ h_{i,t}^{\ell}\geq v_{th},\\ 0,\leavevmode\nobreak\ \leavevmode\nobreak\ otherwise,\end{array}\right.$ (4)
@@
where $I_{i,t}^{\ell}$ represents the input current of the $i$-th neuron in the $l$-th layer at timestep $t$, $u_{i,t}^{\ell}$ represents the membrane potential post-spike firing and reset, and $h_{i,t}^{\ell}$ represents the integrated membrane potential prior to firing. The threshold for membrane potential is denoted as $v_{th}$, and $o_{i,t}^{\ell}$ represents the output following spike firing. The function $g(\cdot)$ in Equation (1) corresponds to different types of spiking neurons, while $\Theta(\cdot)$ in Equation (4) represents the Heaviside step function. In Equation (5), $v_{r}$ stands for the reset potential. This equation elucidates a hard reset scenario in a spiking neuron, where the membrane potential is reset from $u_{i,t}^{\ell}$ to $v_{r}$ upon firing. In surrogate training, the hard reset method is favored over the soft reset approach for enhanced performance [8].
By extending the Equations (1) to (5) to Leaky Integrate-and-Fire (LIF) neuron [9], i.e. $h_{i,t}^{\ell}=\tau_{m}u_{i,t-1}^{\ell}+I_{t}$, we obtain the Equation (6) as:
$u_{i,t}^{\ell}=(\tau_{m}u_{i,t-1}^{\ell}+I_{i,t}^{\ell})(1-\Theta(\tau_{m}u_{i,t-1}^{\ell}+I_{i,t}^{\ell}-v_{th})),$ (6)
where $\tau_{m}$ represents the membrane potential constant, input decay is disregarded, and $v_{r}$ is set to 0 in this work, the iterative update of a spiking neuron adheres to a sequential, incremental process. At time step $t$, the output is contingent on the preceding time step $t-1$. The computational time complexity of the neuron at this stage is $O(T)$, with $T$ denoting the total number of time steps. Given the assumption of independence of membrane potential updates from the previous time point’s output, spiking neuron updates can be architected for parallel processing.
Apparently, when $t=0$, the previous time point’s membrane potential of the neuron is noted as 0, i.e. $u_{i,-1}^{\ell}=0$. Consequently, the computation for $u_{i,0}^{\ell}$ is exact. Referring to Equation (6), we obtain:
$u_{i,0}^{\ell}=I_{i,0}^{\ell}(1-\Theta(I_{i,0}^{\ell}-v_{th})),$ (7)
@@
$I_{i,t}^{\ell}$ $=\sum_{j=1}^{N_{\ell-1}}w_{i,j}^{\ell-1}o_{j,t-1}^{\ell-1},$ (8)
where $w_{i,j}^{\ell-1}$ represents the synaptic weight between the $j$-th neuron in the preceding layer and the $i$-th neuron in the subsequent layer. $N_{\ell-1}$ denotes the number of neurons in the $(\ell-1)$-th layer.
The accurate $u_{i,0}^{\ell}$ is utilized to rectify the error, denoted as $\mathcal{E}$. The procedure is outlined as follows:
$u_{i,t}^{\ell}$ $=(\tau_{m}\hat{u}_{i,t-1}^{\ell}+I_{i,t}^{\ell})(1-\Theta(\tau_{m}\hat{u}_{i,t-1}^{\ell}+I_{i,t}^{\ell}-v_{th})),$ (9)
$u_{i,t+1}^{\ell}$ $=(\tau_{m}\hat{u}_{i,t}^{\ell}+I_{i,t+1}^{\ell})(1-\Theta(\tau_{m}\hat{u}_{i,t}^{\ell}+I_{i,t+1}^{\ell}-v_{th})),$ (10)
@@ 2.2 Initialization of Estimated Membrane Potential
The output of spiking neurons adheres to the $Bernoulli$ distribution. According to the relevant definition in [10], this is described as:
$o_{i,t}^{\ell}\sim Bernoulli(\mathbb{P}[o_{i,t}^{\ell}=1]),$ (12)
where $\mathbb{P}[o_{i,t}^{\ell}=1]$ represents the probability that $o_{i,t}^{\ell}=1$. In our study, we employed the $sigmoid$ function and the input current $I_{i,t}^{\ell}$ to compute $\mathbb{P}[o_{i,t}^{\ell}=1]$, thereby evaluating the firing probabilities of each neuron. Subsequently, the computed probability is utilized to estimate the membrane potential. The calculation procedures in matrix format are outlined as follows:
$\mathbb{P}[\mathbf{o}=\mathbf{1}]$ $=sigmoid(\mathbf{I}),\leavevmode\nobreak\ \mathbf{I}\in\mathbb{R}^{T\times C\times H\times W}$ (13)
$\mathbf{\hat{u}}$ $=(1-Bernulli(\mathbb{P}[\mathbf{o}=\mathbf{1}]))\odot\mathbf{I}.$ (14)
@@
where bold font denotes variables that are matrices. $T$ represents the time steps of the simulation, while $C$, $H$, and $W$ denote the channel, height, and width of the accumulated frames, respectively. The symbol $\odot$ represents the Hadamard product. The time complexity for computing $\mathbb{P}[\mathbf{o}=\mathbf{1}]$ is theoretically $O(1)$, as is that of the $Bernoulli$ operator. Consequently, the additional time required for this operation is negligible, particularly on graphics processing unit (GPU) devices.
@@ 2.3 Membrane Potential Approximation Loss
$\mathcal{L}_{mem}=\sum_{t=0}^{T-1}\sum_{i=0}^{N_{\ell}-1}\kappa_{i}MSE(\hat{u}_{i,t}^{\ell},{u}_{i,t}^{\ell}),$ (15)
The overall loss is calculated as:
@@
where $\lambda$ represents a weighting coefficient, set to 0.01. $\mathcal{L}_{\text{cls}}$ denotes the classification loss, which is computed employing the methodology outlined in the [13]. The loss function of $\mathcal{L}_{cls}$ is $\frac{1}{T}\sum_{t=0}^{T-1}\mathcal{L}_{ce}[\mathbf{o}_{t},\mathbf{y}],$ where $\mathbf{o}_{t}$, $\mathbf{y}$, and $\mathcal{L}_{ce}$ represent outputs, labels, and cross entropy loss respectively.
@@ 3.1 Experimental Details
$\frac{\partial L}{\partial w_{i,j}^{\ell}}=\sum_{t=1}^{T}\frac{\partial L}{\partial o_{i,t}^{\ell}}\frac{\partial o_{i,t}^{\ell}}{\partial{u}_{i,t}^{\ell}}\frac{\partial{u}_{i,t}^{\ell}}{\partial I_{i,t}^{\ell}}\frac{\partial I_{i,t}^{\ell}}{\partial w_{i,j}^{\ell}},$ (17)
@@
$\frac{\partial o_{i,t}^{\ell}}{\partial{u}_{i,t}^{\ell}}=\frac{1}{\alpha^{2}}max(0,\alpha-|{u}_{i,t}^{\ell}-v_{th}|).$ (18)
here, $\alpha$ is a constant that restricts the range of activated gradients, fixed at 1.0. The hyper-parameter configurations and network architecture details are outlined in Table 1. Additionally, $\tau_{m}$ is set to 0.25, with the threshold represented as a parameter subject to learning.
@@ 3.2 Comparisons With Existing Methods
DVSGesture [15]. We performed extensive classification experiments on the DVSGesture dataset using the 7B-Net architecture. Our method achieved a top-1 accuracy of 97.92% at 16 time steps. This outcome surpassed the SOTA result with 4 fewer time steps.
Table 2: Compare with previous studies on neuromorphic and static datasets. T: time steps. B: learnable blocks. Prefix M signifies the term modified. Results marked with $\ast$ denote self-implemented findings. The threshold parameter for Vanilla LIF and PLIF is established at 1.0 for the N-Caltech 101 dataset.
@@
DVS Gesture KLIF  [4] M-PLIF-Net 12 94.1
@@
MPE-PSN 7B-Net 16,20 97.92,97.22
CIFAR10 DVS GLIF [3] 7B-wideNet 16 78.1
@@
N-Caltech 101 [19]. The efficacy of our method was further confirmed on the N-Caltech 101 dataset, with the performance of other methods reproduced for comparison. The outcomes revealed that our approach achieved the highest accuracy presently attainable on the N-Caltech 101 dataset.
@@ 3.3 Membrane Potential Approximation Analysis
(a) $L_{2}\leavevmode\nobreak\ norm(\hat{\mathbf{u}},\mathbf{u})$ and Acc.
@@
(c) Total Loss and Acc.
Fig. 4: 2D loss contours, as introduced in [21], along with accuracy and loss curves, are presented for the CIFAR10DVS dataset with a time step of 4.
@@ 3.4 Computational Efficiency Analysis
A comparative analysis is conducted to assess the computational speed difference between our proposed neuron and the traditional LIF neuron across various time steps and neuron quantities. As depicted in Fig. 1, our neuron generally outpace vanilla LIF neuron in simulation speed. Furthermore, as shown in Fig. 3 (b), our proposed neuron exhibits a notably higher spike rate across most layers when utilizing probabilistic estimation, contrasting with the dynamic complexity observed in the vanilla LIF neuron. Fig. 4 (a) and 4 (b) illustrate that the loss contours of our proposed neuron are comparable to those of vanilla LIF neurons, with both displaying relatively flat profiles. Fig. 3 (a) and 4 (c) demonstrate that initial uncertainties in membrane potential estimation did impact the network’s fitting performance. However, as the estimation of membrane potential improved, this effect became less pronounced. Consequently, our proposed neuron maintain both computational efficiency and performance integrity.